OpenStudy (anonymous):

Use the gradient to find the directional derivative of the function at P in the direction of Q: f(x,y) = sin(2x)cos(y), P(pi,0),Q(pi/2,pi)

OpenStudy (experimentx):

first of all find PQ

OpenStudy (anonymous):

I need help finding u, I'm not sure if I've done this right: PQ = \[PQ=<-\pi/2,\pi>\] \[u=-\pi/2,\pi/\sqrt{(\pi^{2}/4)+(\pi^2)}=5\pi^2/4\]

OpenStudy (experimentx):

let us suppose PQ = r

OpenStudy (experimentx):

df/dr = grad(f).r' --- where r' is unit vector along r

OpenStudy (experimentx):

PQ = \[i \pi/2 - j \pi\]

OpenStudy (amistre64):

gradF with P applied, dotted with Q right?

OpenStudy (anonymous):

\[D_{u}(\pi,0)=2/\sqrt{5}\]

OpenStudy (anonymous):

@JerJason, the u should be a vector which is not just represented by that only. the u is (1/sqrt(5),2/sqrt(5))

OpenStudy (experimentx):

OP and OQ are two position vectors, PQ is a vector that you get from OQ - OP and pi sqrt(5)/2 is the magnitude of PQ, yo can use this magnitude to find the unit vector ... and if you can find the gradient of the function then you can find the directional derivative by taking dot product of gradient and this newly found unit vecotr

OpenStudy (anonymous):

@experimentX P is just a point on the function and Q is the vector which is parallel to the directional derivative

OpenStudy (amistre64):

gradF at P = <-1,0> unit Q-P = <-sqrt(5)pi^2/4, sqrt(5)pi^2/2> dotted = sqrt(5)/4 pi^2

OpenStudy (amistre64):

is Q a vector? or a point? since notation is not consistent between authors

OpenStudy (experimentx):

@anonymoustwo44 I think .. P and Q are two points. ---> ' P(pi,0),Q(pi/2,pi) ' and P in the direction of Q ---> along PQ

OpenStudy (amistre64):

thats what i interped it as as aswell, the vector from P to Q

OpenStudy (anonymous):

it is stated that we are to the the directional derivative of the function a POINT P in the DIRECTION OF Q. so the (x,y) here is P and the u here is vector Q divided by it's magnitude

OpenStudy (anonymous):

@ experimentX at P in the direction of Q not just P in direction of Q

OpenStudy (amistre64):

i would have to defer to the person who posted the question for clarification

OpenStudy (experimentx):

u = PQ/(pi*sqrt(5)/2)

OpenStudy (anonymous):

Yes P and Q are both two points.

OpenStudy (experimentx):

then use the method as i described above